# MiMiC: A High-Performance Framework for Multiscale Molecular Dynamics Simulations

### v1

@@ IV.1 Upcoming Features
Tinker-HPLagardère et al. (2018); Adjoua et al. (2021) is a powerful, massively MPI parallelized MD package, distributed freely for academic institutions, that supports simulations using standard fixed point-charge and advanced polarizable force fields. One of its core strengths is the ability to run long timescale simulations of very large biomolecular systems using polarizable force fields. This is facilitated by a parallel 3D spatial decomposition tailored for point-dipole polarizable models and the use of efficient iterative and non-iterative polarization solvers. Additionally, thanks to recent expansion to GPUs through OpenACC and CUDA, Tinker-HP performs efficiently on both CPU and GPU architecturesAdjoua et al. (2021).
OpenMMEastman et al. (2012, 2017, 2023) is a free and open source classical MD package, distributed under LGPL version 3.0 and MIT licenses, designed to be simple and easy to use, with a focus on extensibility. It offers many common classical force fields, and users can easily add support for new types of force fields via plugins in a modular and extensible way. Notable plugins include polarizable force fields (AMOEBAPonder et al. (2010) and Drude oscillatorLamoureux and Roux (2003) models) and, more recently, ML potentialsEastman et al. (2023). Moreover, it also provides several custom force classes that make it possible for users to specify arbitrary algebraic expressions for force components. It performs efficiently on different hardware platforms, especially GPU-based architectures, for both single- and double-precision calculations.
V Illustrative Applications
@@
Ion channels represent another class of large membrane-protein systems. A recent studySchackert et al. (2023) employed MiMiC to investigate calcium conduction in the $\alpha$-amino-3-hydroxy-5-methyl-4-isoxazolepropionic acid receptors (AMPARs), neurotransmitter-activated cation channels ubiquitously expressed in vertebrate brains (Fig. 8d). In particular, a Ca2+ ion and its first hydration shell were treated at the BLYP or B3LYP levels giving a QM subsystem of between 22 and 25 atoms depending on the simulation, while all other water molecules, as well as protein and lipids, were described classically for a total of 187 699 atoms. In this study, thanks to the QM/MM partitioning, the large electric field of the protein was modeled, as well as the structure and dynamics of calcium-bound water molecules. Moreover, the study provided data to assess the accuracy of a newly developed force field for Ca2+, an ion that is notoriously difficult to describe at the force field level.
MiMiC-based QM/MM simulations have also been used in drug design studies. A combined force field metadynamics and QM/MM simulation study performed in 2020Capelli et al. (2020) investigated the residence time of iperoxo to the human muscarinic acetylcholine receptor 2 (M2). Iperoxo serves as a known M2 antagonist, and therefore its study yields insight into potential therapeutics aimed at M2 receptors in the heart where they regulate the sinus rhythm of the heartbeat rate. The ligand unbinding rate constant $k_{\mathrm{off}}$, and its inverse, the residence time, are crucial pharmacological parametersAhmad et al. (2022). An accurate estimation of $k_{\mathrm{off}}$ requires a precise calculation of the free energy transition state, and not just of the binding free energies and affinities. The QM/MM model places the ligand, key side chains in the binding pocket, and nearby water molecules in the QM subsystem described at the B3LYP level. The MiMiC-based calculations here point to limitations of the force field in the prediction of this important quantity that is due to the lack of charge polarization at the transition state.
@@ VI Outlook
We presently devote the majority of our efforts towards multiscale simulations capable of exploiting the vast computational resources available on current and upcoming exascale supercomputersRemmel (2022); Bolnykh et al. (2021); Martin et al. (2022); Beck, Carloni, and Asthagiri (2024). MiMiC is prepared for the future not only in terms of scalability, but also, the possibility of running the client programs independently and concurrently on separate resources makes it an ideal framework for modular, heterogeneous HPC architecturesCarpenter et al. (2022); Suarez et al. (2022); Carrasco-Busturia et al. (2024). Specifically, each program can use the type and amount of resources that are optimal for its performance, and the different programs can easily run on different partitions or modules of a supercomputer. Hybrid CPU/GPU technology is already prevalent on most state-of-the-art supercomputers, while modular architectures have been adopted by multiple HPC centers. A prominent example is the first European exascale supercomputer JUPITER that will feature multiple modules, including a general-purpose CPU partition and a high-performance GPU-accelerated partitionjup (2024). Moreover, it is expected that future technologies, such as quantum or neuromorphic computing, will eventually be integrated into traditional HPC architectures. As these novel computing technologies become available, MiMiC is virtually ready to incorporate them into its simulation workflows.
@@
In terms of new features, we aim to extend the range of multiscale techniques to address scientific challenges beyond the scope of current QM/MM methods. One of our development directions is to introduce additional versatility in the QM subsystem treatment by enabling larger, more accurate, and yet efficient simulations through fragment-based methods and multilayered QM/QM/MM models. To model larger systems over longer time scales, we will also consider CG and continuum models (CM), enabling, for example, the implementation of MM/CG, QM/MM/CG, and QM/MM/CM schemes. Another promising way to achieve more efficient calculations is the introduction of ML-based force fieldsUnke et al. (2021); Kocer, Ko, and Behler (2022), which will give access to, e.g., ML/MM simulations. Along the same lines is the development of ML-based schemes where a subsystem is described by an ML model that is retrained on the fly based on, e.g., QM calculations. Among the other desired extensions is the implementation of schemes that allow subsystems to adjust to changes over the course of a simulation, such as adaptive QM/MMMato et al. (2021). Finally, we would like to take MiMiC beyond the scope of biomolecular systems, by implementing necessary features for systems within the material science domain. Here, multiscale simulations pose different challenges, but they also open up numerous opportunities to provide a fundamental understanding of various processes, for which current models would not be applicable.

### v2

@@ V Upcoming Features and Client Programs
Tinker-HPLagardère et al. (2018); Adjoua et al. (2021) is a powerful, massively MPI parallelized MD package, distributed freely for academic institutions, that supports simulations using standard fixed point-charge and advanced polarizable force fields. One of its core strengths is the ability to run long timescale simulations of very large biomolecular systems using polarizable force fields. This is facilitated by a parallel 3D spatial decomposition tailored for point-dipole polarizable models and efficient iterative and non-iterative polarization solvers. Additionally, thanks to recent expansion to GPUs through OpenACC and CUDA, Tinker-HP performs efficiently on both CPU and GPU architecturesAdjoua et al. (2021).
OpenMMEastman et al. (2012, 2017, 2023) is a free and open source classical MD package, distributed under LGPL version 3.0 and MIT licenses, designed to be simple and easy to use, with a focus on extensibility. It offers many common classical force fields, and users can easily add support for new types of force fields via plugins in a modular and extensible way. Notable plugins include polarizable force fields (AMOEBAPonder et al. (2010) and Drude oscillatorLamoureux and Roux (2003) models) and, more recently, ML potentialsEastman et al. (2023). Moreover, it also provides several custom force classes that allow users to specify arbitrary algebraic expressions for force components. It performs efficiently on different hardware platforms, especially GPU-based architectures, for both single- and double-precision calculations.
VI Illustrative Applications
@@
Ion channels represent another class of large membrane-protein systems. A recent studySchackert et al. (2023) employed MiMiC to investigate calcium conduction in the $\alpha$-amino-3-hydroxy-5-methyl-4-isoxazolepropionic acid receptors (AMPARs), neurotransmitter-activated cation channels ubiquitously expressed in vertebrate brains (Fig. 9d). In particular, a Ca2+ ion and its first hydration shell were treated at the BLYP or B3LYP levels, resulting in a QM subsystem of between 22 and 25 atoms depending on the simulation, while all other water molecules, as well as protein and lipids, were described classically for a total of 187 699 atoms. In this study, thanks to the QM/MM partitioning, the large electric field of the protein was modeled, as well as the structure and dynamics of calcium-bound water molecules. Moreover, the study provided data to assess the accuracy of a newly developed force field for Ca2+, an ion that is notoriously difficult to describe at the force field level.
MiMiC-based QM/MM simulations used in drug design studies have pointed to the limitations of classical force field descriptions in the prediction of important pharmacological quantities, such as the ligand unbinding rate constant, $k_{\mathrm{off}}$, and its inverse, the residence time. Accurate estimation of $k_{\mathrm{off}}$ requires a precise calculation of the free energy transition state and not just of binding free energy and affinityAhmad et al. (2022). A study in 2020Capelli et al. (2020) investigated the residence time of iperoxo targeting the human muscarinic acetylcholine receptor 2 (M2), yielding insight into potential therapeutics aimed at these receptors, which regulate several key biological processes. This study combined force field metadynamics with MiMiC-based QM/MM simulations, in which the QM subsystem consisted of the ligand, key amino acid side chains, and nearby water molecules described at the B3LYP level.
@@ VII Outlook
We presently devote the majority of our efforts towards multiscale simulations capable of exploiting the vast computational resources available on current and upcoming exascale supercomputersRemmel (2022); Bolnykh et al. (2021); Martin et al. (2022); Beck, Carloni, and Asthagiri (2024). The MiMiC framework is prepared for the future, not only in terms of scalability and extensibility, but also in its capability to run client programs independently and concurrently on separate resources, making it ideal for modular, heterogeneous HPC architecturesCarpenter et al. (2022); Suarez et al. (2022); Carrasco-Busturia et al. (2024). Specifically, each program can use the type and amount of resources that are optimal for its performance, and the different programs can easily run on different partitions or modules of a supercomputer. Hybrid CPU/GPU technology is already prevalent on most state-of-the-art supercomputers, while modular architectures have been adopted by multiple HPC centers. A prominent example is the first European exascale supercomputer JUPITER that will feature multiple modules, including a general-purpose CPU partition and a high-performance GPU-accelerated partitionjup (2024). Moreover, it is expected that future technologies, such as quantum or neuromorphic computing, will eventually be integrated into traditional HPC architectures. As these novel computing technologies become available, MiMiC is virtually ready to incorporate them into its simulation workflows.
@@
In terms of new features, we aim to extend the range of multiscale techniques to address scientific challenges beyond the scope of current QM/MM methods. One of our development directions is to introduce additional versatility in the QM subsystem treatment by enabling larger, more accurate, and yet efficient simulations through fragment-based methods and multilayered QM/QM/MM models. To model larger systems over longer time scales, we will also consider CG and continuum models (CM), enabling, for example, the implementation of MM/CG, QM/MM/CG, and QM/MM/CM schemes. Another promising way to achieve more efficient calculations is the introduction of ML-based force fieldsUnke et al. (2021); Kocer, Ko, and Behler (2022), which will give access to, e.g., ML/MM simulations. Along the same lines is the development of ML-based schemes where a subsystem is described by an ML model that is retrained on the fly based on, e.g., QM calculations. Among the other desired extensions is the implementation of schemes that allow subsystems to adjust to changes over the course of a simulation, such as adaptive QM/MMMato et al. (2021). Finally, we aim to extend MiMiC beyond the scope of biomolecular systems by implementing the necessary features for systems within the material science domain. Here, multiscale simulations pose different challenges, but they also open up numerous opportunities to provide a fundamental understanding of various processes for which current models would not be applicable.